OpenStudy (anonymous):

Angelo Larragu's savings account has a principal of 800. It earns 6 percent interest compounded quarterly. What is the amount in the account at the end of the second quarter? How much is the compound interest?

OpenStudy (kropot72):

The formula to use is as follows: \[A=P(1+\frac{0.06}{4})^{2}\] where A is the amount at the end of the second quarter and P is the principal. When you have found A, the amount at the end of the second quarter, subtract the principal from A to find the compound interest.

OpenStudy (anonymous):

I'm sorry,but I still do not understand.

OpenStudy (anonymous):

http://qrc.depaul.edu/StudyGuide2009/Notes/Savings%20Accounts/Compound%20Interest.htm \[\LARGE A=P(1+\frac{r}{n})^{nt}\] P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year For you: P = 800 r = 0.06 t = 0.5 n = 4 \[\LARGE A=800(1+\frac{0.06}{4})^{4*0.5} = 800(1+\frac{0.06}{4})^{2}\]

OpenStudy (anonymous):

A = the amount in the account at the end of the second quarter. A-P = compound interest